OpenStudy (anonymous):

1. The weight of a certain type of brick has an expectation of 1.12kg with a standard deviation of 0.03kg. (b) How many bricks need to be selected so that their average weight has a standard deviation of no more than 0.005kg?

OpenStudy (kropot72):

The standard deviation of the sample mean is given by \[\sigma _{\bar{x}}=\frac{\sigma}{\sqrt{n}}\] The maximum permitted value of the standard deviation of the sample mean is 0.005kg. Therefore \[0.005=\frac{0.03}{\sqrt{n}}\ ............(1)\] Now you need to solve equation (1) to find n, the smallest sample size.